Hello, this is a logarithmic problem. I need a step by step explanation of:
16^log4(8). I know the answer is 64, but I need to show my work.

Answer by rothauserc(4718)

log (base 4) 8 = x
:
definition of logarithm gives us
:
8 = 4^x
:
x = 3/2
:
4^(3/2) says take square root of (4^3)
:
therefore 16^(3/2) = square root(16^3) = 64
: